use the following data to create a balance sheet and an income statement place each item in the proper category example (assets liabilities income expense) then calculate the net worth and the net income.
cash $1000
monthly revenue $2000
current credit card balance $1000
value of inventory $1000
monthly rent expense $500
current automobile value $7000
current value of equipment $5000
monthly insurance expense $100
automobile loan balance $4000
monthly wage expense $500
Answer #1
Assets Liability
Cash 1000 credit card balance 1000
Value of inventory 1000 automobile loan balance 4000
Current automobile value 7000
Current value of equipment 5000
Total 14000 total 5000

Net worth = 14000-5000 = 9000$

Income Expenses
Monthly revenue 2000 monthly rent 500
Wage 500
insurance 100
Total 2000 total 1100

Net income = 2000-1100 = 900$